The average rainfall in Waterloo has increased at an annual rate of 2%. If the total rain fall was 856mm in 2002. What is the expected rainfall in 2005? (mm)
Answer by Theo(13342)

856 * 1.02 ^ (2005 - 2002) = 856 * 1.02 ^ 3 = 908.394048.
the expected rainfall in 2005 is 908.394048mm.